OpenStudy (joannablackwelder):

If a 9 cm dia. ornament is hanging from a tree and the reflection in the ornament is 1/2 the size, how far away is the person viewing it?

OpenStudy (joannablackwelder):

@IrishBoy123

OpenStudy (joannablackwelder):

@Michele_Laino

OpenStudy (joannablackwelder):

Oh, I think this ends up being just a simple 1/f= 1/do + 1/di problem, right?

OpenStudy (irishboy123):

yes it does

OpenStudy (irishboy123):

so f = -2.25, i think.......

OpenStudy (michele_laino):

we have to consider the formula of magnification

OpenStudy (joannablackwelder):

Right, thanks so much!

OpenStudy (michele_laino):

for example, if we write this: \[\frac{1}{p} + \frac{1}{q} = \frac{1}{f}\] where \(p\) is the distance of the object and \(q\) is the distance of the image, then the magnification \(G\) is given by the subsequent formula: \[G = \frac{q}{p}\]

OpenStudy (michele_laino):

In your case, since the magnification is equal to 2, then we can write: \[G = \frac{q}{p} = 2 \Rightarrow q = 2p\]

OpenStudy (michele_laino):

so we get: \[\frac{1}{p} + \frac{1}{q} = \frac{1}{f} \Rightarrow \frac{1}{p} + \frac{1}{{2p}} = \frac{4}{D}\] where \(D= 9\) cm, please solve that equation for \(p\)

OpenStudy (joannablackwelder):

Hm, where did the 4 come from?

OpenStudy (irishboy123):

half of half of diameter is focal length...

OpenStudy (michele_laino):

correct! @IrishBoy123 we can write this: \[f = \frac{R}{2} = \frac{{D/2}}{2} = \frac{D}{4} \Rightarrow \frac{1}{f} = \frac{4}{D}\]